Circular No. 5611

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 QB_{1}

D. Jewitt, University of Hawaii; and J. Luu, University of California at Berkeley, report the discovery of a very faint object with very slow (3′′/hour) retrograde near-opposition motion, detected in CCD images obtained with the University of Hawaii’s 2.2-m telescope at Mauna Kea. The object appears stellar in 0′′.8 seeing, with an apparent Mould magnitude R = 22.8 ± 0.2 measured in a 1′′.5-radius aperture and a broadband color index V − R = +0.7 ± 0.2.

1992 UT  α_{2000}  δ_{2000}
Aug. 30.45568  0^{h}01^{m}12.^{s}79  +0^{o}08′50′′.7
30.59817  0 01 12.19  +0 08 46.9
31.52047  0 01 08.37  +0 08 22.7
31.61982  0 01 07.95  +0 08 19.9
Sept. 1.35448  0 01 04.90  +0 08 00.6
1.62225  0 01 03.76  +0 07 53.3

Computations by the undersigned indicate that 1992 QB_{1} is currently between 37 and 59 AU from the earth but that the orbit (except for the nodal longitude) is completely indeterminate. Some solutions are compatible with membership in the supposed “Kuiper Belt”, but the object could also be a comet in a near-parabolic orbit. The particular solution below is the direct circle (but a retrograde circle some 15 AU larger in radius also fits); Jewitt and Luu note that a cometlike albedo of 4 percent then implies a diameter of 200 km and that the red color suggests a surface composition rich in organics. Further precise astrometry during the late-September dark run should eliminate some possibilities, but a satisfactory definition of the orbit will clearly require follow-up through the end of the year. The object’s phase angle reaches a minimum of < 0^{o}.01 around Sept. 22.5 UT.

Epoch = 1992 Aug. 26.0 TT  u = 0.^{o}335
Ω = 359.440 }2000.0
a  = 41.197 AU  i = 2.334
1992TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Sept. 15 0^{h}00.^{m}09 + 0^{o}01.^{′}7 40.200 41.197 172.^{o}5 0.^{o}2 23.4
25 23 59.33 − 0 03.1 40.195 41.197 177.5 0.1 23.4
Oct. 5 23 58.58 − 0 07.9 40.220 41.197 167.5 0.3 23.5
15 23 57.87 − 0 12.5 40.275 41.197 157.4 0.5 23.5

1992 September 14 Brian G. Marsden